a bag contains a selection of multilink cubes. there are 3 red, 5 yellow, 2 blue and 6 green cubes. a cube is picked at random. what is the probability that a blue or yellow cube is NOT chosen?
"my answer is 1/2 am i right"

Answer by ewatrrr(24785)

16cubes: 3R,5Y,2B,6G
one picked
not B 0r Y then R 0r G
P(R 0r G) = 3/16 + 6/16 = 9/16